AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 06:33:41 2025
Author
*
:
Subject
*
:
<p>15:27</p> <p>Run 40 started cooled beam on target</p> <p>Detectors IN (left arm pot =36.5 mm, right arm pot = 70.49 mm). Stationary. Same as run 35</p> <p>Electron cooler voltage 751.095 this corresponds to 1.338 MeV/u. This should be on the left of the peak of the resonance.</p> <p>Target density is 4e11 at/cm2, ring pressure looks OK YR09 1.43e-10 mbar.</p> <p>Horizontal beam position is -7.5 mm, vertical beam position is 0.0 mm. Same as run 35.</p> <p>Beam is in the ring for 4 seconds and total cycle time between injections is 8 seconds. Notice this is different than run 35.</p> <p>Statistics, merger statistics, histograms all zero'd</p> <p>15:39</p> <p>System wide checks ok. Merger ok. Temperatures ok.</p> <p>Attatchment 1: statistics</p> <p>Attatchment 2: Temperatures</p> <p>Attatchment 3-6: Waveforms. Layout 2 - (attach 4) had a strnageish waveform for aida11, stays after reload.It was the HEC firing so did an ASIC control and Zero'd histograms and all is okay.</p> <p>Attatchment 7: Leakage current</p> <p>17:04</p> <p>System wide checks ok. Merger ok. Temperatures ok. The beam is on target and everything runs smoothly.</p> <p>Attatchment 8: statistics</p> <p>Attatchment 9: Temperatures</p> <p>Attatchment 10-13: Waveforms. Everything looks ok.</p> <p>Attatchment 14: Leakage current</p> <p>18:55</p> <p>System wide checks ok. Merger ok. Temperatures ok. The beam is on target and everything runs smoothly.</p> <p>Beam transformer shows a small decrease in beam intensity, not drastic but small slow decrease, see attachment 22. Target density 4e11. Vacuum pressue ok.</p> <p>Attatchment 15: Statistics</p> <p>Attatchment 16: Temperatures</p> <p>Attatchment 17-20: Waveforms</p> <p>Attatchment 21: Leakage current</p> <p>19:28</p> <p>Run40 - Beam dropped intesntiy at 19:17, moved detectors out, called gleb. Beam intensity started increasing again - not sure whats happenig at present (see image 23). </p> <p>20:26</p> <p>Around 19:50 beam decreased again and stayed down for a while. Beam not stable. Gleb was called, target turned off, beam restabilied - around 2.5e7 ish. Unsure if it will remain stable during the night (image 24).</p> <p>20:35</p> <p>detectors moved back in. Target back in, stable at 4e11.</p> <p>Potentiometer Left: 36.54 mm</p> <p>Potentiometer Right: 70.49 mm</p> <p>Beam stable at the moment (image 25).</p> <p> </p> <p>System checks 20:40</p> <p> <br /> Base Current Difference<br /> aida03 fault 0x0 : 0x1 : 1 <br /> aida04 fault 0x0 : 0x1 : 1 <br /> aida05 fault 0x0 : 0x1 : 1 <br /> aida06 fault 0x0 : 0x1 : 1 <br /> aida07 fault 0x0 : 0x1 : 1 <br /> aida08 fault 0xe : 0xf : 1 <br /> aida09 fault 0xb : 0xc : 1 <br /> aida10 fault 0x0 : 0x1 : 1 <br /> aida13 fault 0x0 : 0x1 : 1 <br /> aida14 fault 0x0 : 0x1 : 1 <br /> FPGA Timestamp error counter test result: Passed 6, Failed 10<br /> If any of these counts are reported as in error<br /> The ASIC readout system has detected a timeslip.<br /> That is the timestamp read from the time FIFO is not younger than the last</p> <p>System wide checks ok except the FPGA timestamps - see error above - collected them and it went away. Merger ok. Temperatures ok. The beam is on target and everything runs smoothly for now.</p> <p>Attatchment 26: Temperatures</p> <p>Attatchment 27-30: Waveforms</p> <p>Attatchment 31: Leakage current</p> <p> </p> <p>21:50</p> <p>FEE temperatures OK</p> <p>System wide checks OK</p> <p>Statistics and spectrum browser rate look OK</p> <p>Merger and tape server OK</p> <p>Monitoring code rates are consistent with beam on target</p> <p>X2go seems to not be loading full screen for me - may be my laptop misbehaving?</p> <p>Can see ring intensity has been stable for the last hour, target density is 4.33e11 and YR09 pressure is 1-2e-10 mbar.</p> <p> </p> <p>23:22</p> <p>Target density 4.3e11, beam intensity stable.</p> <p>Monitoring code rates are consistent with beam on target</p> <p>Fee temps all OK</p> <p>System checks OK</p> <p>Statistics and FEErates OK.</p> <p>Merger and tape server OK</p> <p> </p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown